### v1

# Nonlinearity-induced Thouless pumping of solitons

Abstract
It has recently been theoretically predicted and experimentally observed that a soliton resulting from nonlinearity can be pumped across an integer or fractional number of unit cells as a system parameter is slowly varied over a pump period. Nonlinear Thouless pumping is now understood as the flow of instantaneous Wannier functions, ruling out the possibility of pumping a soliton across a nonzero number of unit cells over one cycle when a corresponding Wannier function does not exhibit any flow, i.e., when the corresponding Bloch band that the soliton bifurcates from is topologically trivial. Here we surprisingly find an anomalous nonlinear Thouless pump where the displacement of a soliton over one cycle is twice the Chern number of the Bloch band from which the soliton bifurcates. We show that the breakdown of the correspondence between the displacement of a soliton and the Chern number arises from the emergence of a new branch of stable soliton solutions mainly consisting of two neighboring instantaneous Wannier functions. Furthermore, we find a nonlinearity-induced integer quantized Thouless pump, allowing a soliton to travel across one unit cell during a pump period, even when the corresponding band is topologically trivial. Our results open the door to studying nonlinearity-induced Thouless pumping of solitons.
Thouless pumping, which plays an important role in understanding the quantum Hall effect, is a phenomenon where quantized transport arises through the slow periodic variation of a Hamiltonian parameter [1]. The transport requires the complete filling of an energy band so that it is dictated by the Chern number of the occupied band with the varying parameter playing the role of an additional momentum. So far, Thouless pumps have been experimentally observed in photonic systems [2, 3, 4], cold atom systems [5, 6, 7, 8, 9] and other systems [10, 11, 12].
Nonlinearity is a natural feature in various systems, including photonic systems [13, 14, 15, 16, 17, 18, 19] and Bose-Einstein condensates (BECs) [20, 21, 22, 23, 24, 25]. It leads to many intriguing phenomena, such as solitons, which are wave packets that travel without changing their shape. Nonlinearity also gives rise to topology-related phenomena, such as topological bulk [26, 27] and edge solitons [28, 29, 30, 31], as well as nonlinearity-induced topological insulators [32, 33]. In particular, it has recently been found that a soliton arising from nonlinearity can exhibit quantized transport when an underlying Hamiltonian parameter is slowly tuned periodically (called nonlinear Thouless pumping) [34, 35, 36, 37, 38, 39]. Later, a quantized fractional Thouless pumping of solitons is found by considering a soliton bifurcating from multiple bands [40]. The integer and fractional quantized Thouless pumping are now understood as the flow of the instantaneous one-band and maximally localized multi-band Wannier functions [35, 36, 37, 40], respectively. For the former, the displacement of a soliton is determined by the Chern number of the Bloch band of the linear Hamiltonian of the system [35, 36, 37]. If this mechanism always worked, we would not observe the occurrence that a soliton’s displacement is larger than this Chern number. In other words, a soliton would not undergo transport when a linear Hamiltonian is topologically trivial.
Here we surprisingly find an anomalous nonlinear Thouless pump that exhibits the pumping of a soliton across two unit cells over one cycle, while the Chern number of the band, from which the soliton bifurcates, is $C=-1$. Our result is thus beyond the previous nonlinear Thouless pump where the displacement is identical to the Chern number. It cannot be attributed to the renormalization of a system parameter resulting from nonlinearity, as variations of parameters only permit the Chern number to change from $-1$ to $1$, rather than to $2$ in our system. We show that the phenomenon arises due to the emergence of a new branch of stable instantaneous soliton solutions formed by a superposition of two neighboring instantaneous Wannier functions. These soliton solutions enable the displacement to be two times the Chern number. Leveraging the discovery of the new nonlinear Thouless pump, we construct an alternative nonlinear model, demonstrating the occurrence of quantized Thouless pumping even when the Chern number of the band of the linear Hamiltonian vanishes (in other words, the linear Thouless pumping does not happen in this system over one cycle). To further characterize these anomalous phenomena, we study a spatial supercell containing multiple unit cells to ensure the well localization of a soliton within the supercell. Subsequently, we calculate the Chern number of a modulated Hamiltonian involving effects of the soliton solution, finding it to be identical to the displacement. Finally, we show the emergence of anomalous nonlinear Thouless pumping in a continuous model, which can be experimentally realized in cold atom systems.
Nonlinear model.— To illustrate our findings, we start by investigating the following dimensionless discrete nonlinear Schrödinger equation:
$i\frac{\partial}{\partial t}\psi_{\sigma x}=$ $\sum_{\sigma^{\prime}x^{\prime}}H_{\sigma x,\sigma^{\prime}x^{\prime}}^{\textrm{lin}}(\theta)\psi_{\sigma^{\prime}x^{\prime}}$ (1)
$+\left(g|\psi_{\sigma x}|^{2}+g_{12}|\psi_{\overline{\sigma}x}|^{2}\right)\psi_{\sigma x},$
where $\psi_{\sigma x}$ is the wavefunction of the $\sigma$th component ($\sigma=1,2$) in the $x$th unit cell at time $t$ ($t$ is the propagation distance in photonic systems), $H^{\textrm{lin}}$ is the linear tight-binding Hamiltonian that depends on a real system parameter $\theta$, $g$ and $g_{12}$ are nonlinear coefficients describing the intraspecies and interspecies interactions in ultracold atomic gases, respectively [$\overline{\sigma}=(\sigma\text{ mod }2)+1$ labels the other component]. The norm of the wavefunction $N=\sum_{\sigma x}|\psi_{\sigma x}|^{2}$ is preserved during the evolution in this equation. To show that the displacement of a soliton can be twice the Chern number of a Bloch band, we choose a linear Hamiltonian which reads in momentum space
$H^{\textrm{lin}}(k)=(m_{z}+J_{1}\cos k)\sigma_{z}+J_{1}^{\prime}\sin k\sigma_{y}+J_{2}\sigma_{x},$ (2)
where $\sigma_{\nu}$ ($\nu=x,y,z$) are Pauli matrices, and $m_{z}=m_{0}+\cos\theta$ with $m_{0}$ being a real system parameter. We set $J_{1}=J_{1}^{\prime}=1$ and $J_{2}=\sin\theta$. The parameter $\theta$ is slowly varied as time evolves, realizing a periodic change of the linear Hamiltonian. Clearly, this model realizes the Chern band [41] if we view $\theta$ as the other momentum $k_{y}$ besides $k$. The Chern number of the lowest band with respect to $k$ and $\theta$ is $C=1$ when $-2<m_{0}<0$, $C=-1$ when $20<m_{0}<$ 2, and $C=0$ otherwise. In the following study concerning a soliton’s displacement of $-2$, we take $m_{0}=1$. For linear Thouless pumping, particles are required to fill the entire band so that the cloud of particles will travel across the lattice distance identical to the Chern number as we slowly tune $\theta$ from 00 to $2\pi$ [1]. Note that this tight-binding model can be experimentally implemented in ultracold atomic gases as proposed in Ref. [42] (see the following discussion).
Figure 1: Anomalous nonlinear Thouless pumping. a, One-cycle evolution of the density distribution $|\psi_{x}|^{2}=\sum_{\sigma}|\psi_{\sigma x}|^{2}$ of instantaneous solitons which bifurcate from the lowest band when $g_{12}=0$. b, The trajectory of center-of-mass positions of instantaneous solitons when $g_{12}=0$ (red line), $g_{12}=1$ (blue line) and the corresponding Wannier functions (black line) as a system parameter $\theta$ varies. Here, we set $N=1.45$ and $g=1$.
Nonlinear evolution and mechanism.— With nonlinearity, spatially localized solutions known as solitons can emerge due to the balance between nonlinearity and dispersion [17, 19, 22, 23]. Remarkably, previous studies have demonstrated that in a nonlinear Aubry-André-Harper model, a soliton can be transported across an integer number of unit cells per cycle, which is dictated by the Chern number of the underlying Bloch band from which the soliton bifurcates [34]. The phenomenon is now understood as the motion of a soliton following the center-of-mass positions (referred to as the Wannier center) of Wannier functions of the corresponding Bloch bands [34, 35, 36, 37, 38, 40]. In our case, given that the Chern number of the lowest band is $C=-1$, we expect that the displacement of a soliton per cycle should be $-1$ (in units of the length of a unit cell, which we will take hereafter) based on the theory. However, contrary to our expectations, we find that the actual displacement is $-2$, deviating from the value predicted by the Chern number.
To demonstrate our findings, we calculate stable instantaneous soliton solutions that bifurcate from the lowest band with the Chern number $C=-1$ for the instantaneous nonlinear Hamiltonian at each $\theta$ using the Newton’s method. This approach is justified because, in the adiabatic limit—where $\theta$ is varied very slowly—the time evolution of a soliton follows the instantaneous eigenstate of the instantaneous nonlinear Hamiltonian at time $t$ [43, 44, 45, 46, 47]. We have directly computed the time evolution based on Eq. (1) when $\theta$ is varied sufficiently slowly and found that the results closely resemble the instantaneous solutions. Additionally, we have performed a stability analysis to confirm that the solved nonlinear instantaneous soliton solutions remain stable at each $\theta$ (see also Supplementary Information Section S-1).
Figure 1b illustrates that when $g_{12}=1$, the soliton travels across the lattices by a distance of $-1$ over one period, which is identical to the Chern number of the lowest band. The result is consistent with the previous findings [34, 35, 36, 37]. Suprisingly, we observe that when $g_{12}=0$, the soliton’s displacement becomes $-2$, twice the Chern number of the lowest band, as shown in Fig. 1. We find that the transition between the two displacements occurs near $g_{12}=0.2$ when $N=1.45$. The phenomenon cannot be attributed to the renormalization of the system parameter $m_{0}$ resulting from nonlinearity, as variations in $m_{0}$ only enable the Chern number to change from $-1$ to $1$, rather than to $2$.
Figure 2: Mechanism of anomalous nonlinear pumping. a-d, Center-of-mass positions of stable instantaneous solitons (black and red lines) and Wannier functions (dashed grey lines) with respect to a system parameter $\theta$ over one cycle for different $g_{12}$. The anomalous nonlinear pumping arises due to the appearance of a new branch of soliton solutions around $\theta=\pi$ plotted as red lines in b. Here, we set $N=1.45$ and $g=1$.
To explain the phenomenon, we plot in Fig. 2 the center-of-mass position of stable nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as a function of $\theta$ for different values of $g_{12}$. Notably, when $g_{12}=0.3$, the center-of-mass positions of a soliton and the Wannier centers exhibit apparent differences for most values of $\theta$, but they coincide at $\theta=0$ and $\theta=\pi$. This results in a displacement of $-1$ over each period, following the trajectory indicated by the black lines. In fact, at $\theta=\pi$, the soliton solution is exactly identical to the Wannier function multiplied by $\sqrt{N}$ when $m_{0}=1$, as will be further elucidated below. However, as $g_{12}$ is decreased, a stable soliton solution emerges at $\theta=\pi$, localized around $x=n$ with $n$ being an integer, differing from the Wannier center, which is located at $n+1/2$. With a further decrease in $g_{12}$, we observe that the new branch of solutions becomes longer, eventually connecting to the original soliton solutions (represented by black lines). Consequently, when $g_{12}<0.2$, a soliton becomes displaced by $-2$ unit cells during adiabatic evolution, following the trajectory indicated by the black lines in Fig. 2c,d.
To elucidate the emergence of the nonlinear solution at $\theta=\pi$ localized around $x=n$, we transform the linear Hamiltonian to the Su-Schrieffer-Heeger (SSH) model, expressed as, $H^{\textrm{lin}}_{\textrm{SSH}}=UH^{\textrm{lin}}U^{\dagger}=(m_{z}+\cos k)\sigma_{x}+\sin k\sigma_{y}$ with $U=e^{-i\pi\sigma_{y}/4}$. When $m_{0}=1$ so that $m_{z}=0$, the model in real space contains only intercell hopping. In this scenario, we show that the SSH model’s Wannier function multiplied by $\sqrt{N}$ is a solution to the transformed nonlinear equation (see Supplementary Information Section S-2). However, we find that these solutions become unstable when $g_{12}$ is small.
For solutions at $\theta=\pi$ that localize between two adjacent Wannier functions, it is likely that these solutions represent a superposition of these Wannier functions. Indeed, we find that these solutions mainly consist of equal superpositions of two neighboring Wannier functions of the lowest band (see Supplementary Information Section S-2). We note that similar solutions have also been found in Ref. [35], but these solutions are unstable in their case. In contrast, although our solutions are unstable for large $g_{12}$, they become stable for smaller values of $g_{12}$, facilitating the anomalous transport of a soliton. While we focus on the specific case of $m_{0}=1$ for simplicity of analysis, we have confirmed that anomalous nonlinear Thouless pumping occurs for other values of $m_{0}$ as well.
Nonlinearity-induced Thouless pumping.— We have demonstrated the breakdown of the correspondence between displacement and the Chern number for a nonlinear Thouless pump, which naturally leads to a question whether a soliton can be pumped when the Chern number of the linear band is zero. Our answer is affirmative. To illustrate this, we again consider the linear Hamiltonian in Eq. (2). Now, we vary the system parameters $m_{z}$, $J_{1}=J_{1}^{\prime}$ and $J_{2}$ with respect to $\theta$, as depicted in Fig. 3(a). For simplicity, we impose the conditions that $m_{z}(\theta)=m_{z}(2\pi-\theta)$, $J_{1}(\theta)=J_{1}(2\pi-\theta)$ and $J_{2}(\theta)=J_{2}(2\pi-\theta)$, ensuring that these parameters are symmetric about $\theta=\pi$. Consequently, the Hamiltonian $H^{\textrm{lin}}(k)$ also exhibits symmetry with respect to $\theta$, i.e., $H^{\textrm{lin}}(k,\theta)=H^{\textrm{lin}}(k,2\pi-\theta)$. The immediate result is that the Berry phase of the $n$th ($n=1,2$) band of $H^{\textrm{lin}}(k,\theta)$ satisfies $\gamma_{n}(\theta)=\gamma_{n}(2\pi-\theta)$ [where $\gamma_{n}(\theta)$ denotes the Berry phase following the path along $k$ for the $n$th band at a fixed $\theta$], leading to a Chern number of zero for each band. Therefore, the linear Thouless pumping cannot occur.
To achieve nonlinearity-induced Thouless pumping of solitons, we simultaneously adjust $g_{12}$, as shown in Fig. 3b (the interaction can be controlled via Feshbach resonances in ultracold atomic gases [24]). Specifically, we initially reduce $g_{12}$ to zero gradually, which facilitates the soliton’s travel over a longer distance. Subsequently, we slowly increase $g_{12}$ back to its original value to prevent the soliton from returning to its initial position. In Fig. 3c, we plot the density distribution of nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as $\theta$ varies from 00 to $2\pi$, remarkably demonstrating that the soliton is displaced by $-1$ unit cell. This effect is further evidenced in Fig. 3d, which depicts the evolution of center-of-mass positions of instantaneous soliton solutions as $\theta$ changes. This observation starkly contrasts with the behavior of the Wannier center (or the Berry phase), which returns to its starting value after one complete cycle. Hence, we successfully achieve Thouless pumping of a soliton through the influence of nonlinearity.
Figure 3: Nonlinearity-induced Thouless pumping. a,b, Illustration of how the parameters $m_{z}(\theta)$, $J_{1}(\theta)$, $J_{2}(\theta)$ and $g_{12}(\theta)$ are varied with respect to $\theta$ in order to induce a nonlinear pumping for a topologically trivial band. c, One-cycle evolution of the density distribution $|\psi_{x}|^{2}$ of instantaneous solitons bifurcating from the lowest band. d, The evolution of center-of-mass positions of instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one cycle. Here, we set $N=1.25$ and $g=1$.
To further characterize the topological property of anomalous nonlinear Thouless pumping, we calculate the Chern number of a modulated linear Hamiltonian in a supercell consisting of $L$ unit cells incorporating the effects of a soliton [34] (see also Supplementary Information Section S-3). We choose an $L$ (here we take $L=40$) such that the soliton is well localized within the supercell. As we vary $\theta$ from 00 to $2L\pi$, the soliton is pumped across $Ld_{p}$ unit cells where $d_{p}$ is the displacement over each cycle of $\theta$ from 00 to $2\pi$. When the soliton travels outside the supercell, we shift it by $L$ back into the supercell. By involving the effects of the soliton, we define a new linear Hamiltonian $H^{\textrm{sc}}(\theta)$ as
$H^{\textrm{sc}}_{\sigma x,\sigma^{\prime}x^{\prime}}:=H^{\textrm{lin}}_{\sigma x,\sigma^{\prime}x^{\prime}}+(g|\psi_{\sigma x}^{\textrm{s}}|^{2}+g_{12}|\psi_{\overline{\sigma}x}^{\textrm{s}}|^{2})\delta_{xx^{\prime}}\delta_{\sigma\sigma^{\prime}},$ (3)
where $\psi_{\sigma x}^{\textrm{s}}$ is the soliton solution we consider at a fixed $\theta$. Under periodic boundary conditions, $\psi_{\sigma x}^{\textrm{s}}$ is clearly an eigenstate of $H^{\textrm{sc}}$. By applying twisted boundary condition for the Hamiltonian $H^{\textrm{sc}}$ with a twisted phase $k_{\textrm{sc}}$ at each $\theta$, we calculate the Chern number $C_{\textrm{sc}}$ of the correponding band of $H^{\textrm{sc}}$ with respect to $\theta^{\prime}=\theta/L$ and $k_{\textrm{sc}}$. We find that $C_{\textrm{sc}}=d_{p}$ when $d_{p}=0,1,2$ in our cases. The results are self-consistent as the soliton in a supercell plays the role of the Wannier function of $H^{\textrm{sc}}$ consisting of many supercells (our numerical results also support this viewpoint). As a result, the soliton will travel across $C_{\textrm{sc}}$ supercells after a period of $\theta^{\prime}$ so that the displacement over one cycle of $\theta$ is $C_{\textrm{sc}}$. The anomalous nonlinear Thouless pumping is thus incorporated into the framework of the supercell theory.
Continuous model.— To implement the tight-binding model in ultracold atomic gases, we consider the following continuous model:
$H_{\textrm{c}}^{\textrm{lin}}=H_{0}+h_{z}\sigma_{z}+V_{\textrm{so}}\sigma_{x},$ (4)
where the spin is encoded in two hyperfine states of an atom. Note that similar models have been proposed for realizing linear Thouless pumping in both two-dimensional and three-dimensional cold atom systems [42]. In this Hamiltonian, $H_{0}=\frac{p_{x}^{2}}{2m}-V_{x}\cos^{2}(k_{R}x)$, where $p_{x}=-i\hbar\partial_{x}$ is the momentum operator, $m$ is the mass of atoms, and $V_{x}$ is the strength of optical lattices generated by lasers with wavevector $k_{R}$. In addition, $h_{z}$ denotes the strength of the Zeeman field correponding to the detuning of the Raman lasers, and $V_{\textrm{so}}=V_{s}\sin(k_{R}x)+V_{c}\cos(k_{R}x)\sin(\theta)$. This potential can be implemented using two pairs of Raman lasers, as described in Ref. [42, 48, 49, 50]. One pair (the other) contains a laser beam with the Rabi frequency proportional to $\sin(k_{R}x)$ [$\cos(k_{R}x)$] and another beam proportional to $e^{ik_{R}y}$. Given that our system is effectively one-dimensional due to a strong confinement along $y$ and $z$, the effects of the phase $e^{ik_{R}y}$ can be negligible. We can map the continuous model to a tight-binding model based on the Wannier functions $\{W_{j}\}$ of the lowest Bloch band of the Hamiltonian $H_{0}$, yielding precisely the Hamiltonian in Eq. (2) with $J_{1}=2t_{1}$, $J_{1}^{\prime}=-2t_{s}$, $J_{2}=t_{c}\sin\theta$ and $m_{z}=h_{z}$ (see Supplementary Information Section S-4 for their definition). Thus, we expect the emergence of anomalous nonlinear Thouless pumping in the continuous model in the presence of interactions described by nonlinear terms.
For the linear Thouless pumping of this model, we prove that the average displacement per particle at time $t$ is $x_{c}=a[\gamma(\theta(t))-\gamma(\theta(t=0))]/(2\pi)$ where $\gamma(\theta(t))$ is the Berry phase following the path along $k$ at a fixed $\theta$ and $a=\pi/k_{R}$, although the length of a unit cell of $H_{c}^{\textrm{lin}}$ is $2a$ due to the presence of $V_{\textrm{so}}$ (see Supplementary Information Section S-5 for proof). Here we require that $\gamma(\theta(t))$ changes continuously from $\theta(0)$ to $\theta(t)$. In fact, we find that the linear eigenstates can be characterized by the momentum $k\in[0,2\pi/a]$. In our study, we set $V_{x}=4E_{R}$, $V_{s}=1.04E_{R}$ and $V_{c}=0.21E_{R}$, where $E_{R}=\hbar^{2}k_{R}^{2}/(2m)$ is the recoil energy. The parameter $h_{z}$ is varied as $h_{z}=2t_{x}(1+\cos\theta)$ with $t_{x}=0.0855E_{R}$. Under these conditions, we find that the Chern number of the first and second bands of the continuous model are $C=1$ and $C=-1$, respectively. As a result, the average displacement for the linear Thouless pumping over one cycle is $a$ and $-a$, respectively (see Fig. 4b).
Figure 4: Anomalous nonlinear pumping in the continuous model. a, One-cycle evolution of the density distribution $|\psi(x)|^{2}=\sum_{\sigma}|\psi_{\sigma}(x)|^{2}$ of the instantaneous solitons bifurcating from the second band when $g_{12}=0$. b, The evolution of center-of-mass positions $x_{c}$ of the instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one period. Here, ${N}=0.2$ and $g=-1$.
We now demonstrate the emergence of anomalous nonlinear pumping in the continuous model in the presence of nonlinearity. The dynamics is governed by the following dimensionless Gross-Pitaevskii equation:
$i\frac{\partial}{\partial t}\psi_{\sigma}(x)=[\tilde{H}_{c}^{\textrm{lin}}\psi(x)]_{\sigma}+g|\psi_{\sigma}(x)|^{2}\psi_{\sigma}(x),$ (5)
[…]
In addition, we show that for negative $g$, a soliton bifurcating from the first band does not exhibit any pumping over one cycle. Although the phenomenon is analogous to a trapped soliton [34], the underlying mechanism differs markedly. The trapped phenomenon arises from multiple band contributions due to strong nonlinearity [34, 36, 40]. In contrast, in our case, strong nonlinearity is not a prerequisite, and the projections of the soliton indicate primary occupation of a single band. This behavior in fact stems from the emergence of a new branch of soliton solutions, which is primarily composed of two adjacent instantaneous Wannier functions (see Supplementary Information Section S-6).
Discussion.— We have demonstrated the emergence of anomalous nonlinear Thouless pumping, wherein the correspondence between the displacement of a soliton and the Chern number of the underlying linear Bloch band breaks down. This anomalous behavior surpasses the understanding of nonlinear Thouless pumping based on the center-of-mass flow of the instantaneous Wannier functions. We show that the breakdown arises due to the emergence of a new branch of soliton solutions primarily composed of two neighboring instantaneous Wannier functions. Furthermore, we illustrate that nonlinearity can induce the Thouless pumping of a soliton, even when the corresponding linear Bloch band is topologically trivial. We note that our results are completely different from the phenomenon in a Bose-Bose atomic mixture in Ref. [37] where a soliton of impurity atoms can undergo a nonzero quantized displacement during a pump cycle, as a result of the dragging from a soliton of majority atoms that bifurcates from a topologically nontrivial band. There, the pumping can also be understood as the center-of-mass flow of the Wannier functions of the corresponding Bloch band of the linear Hamiltonian felt by the majority atoms. In our case, we have only one set of band structures, and the discrepancy between the band topology and quantized motion emerges spontaneously, going beyond the Wannier function description. Given that anomalous nonlinear Thouless pumping can occur in both discrete and continuous models, it can be experimentally observed in either photonic systems or ultracold atomic gases. Additionally, our work inspires the exploration of other interesting topics, such as anomalous fractional Thouless pumping of solitons.
Acknowledgments We thank Y. Zhang and Y.-B. Yang for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604). We also acknowledge the support by center of high performance computing, Tsinghua University.