### v2

# Anomalous quantized nonlinear Thouless pumping

Abstract
It has recently been theoretically predicted and experimentally observed that a soliton resulting from nonlinearity can be pumped across an integer or fractional number of unit cells as a system parameter is slowly varied over a pump period. Nonlinear Thouless pumping is now understood as the flow of instantaneous Wannier functions, ruling out the possibility of pumping a soliton across a nonzero number of unit cells over one cycle when a corresponding Wannier function does not exhibit any flow, i.e., when the corresponding Bloch band that the soliton bifurcates from is topologically trivial. Here we surprisingly find an anomalous nonlinear Thouless pump where the displacement of a soliton over one cycle differs from the Chern number of the Bloch band from which the soliton comes. We develop a general theory showing that this anomalous behavior arises from a transition of a soliton between different Wannier functions by passing through an intersite-soliton (or dipole-soliton) state. Furthermore, we find a nonlinearity-induced integer quantized Thouless pump of a soliton, allowing a soliton to travel across one unit cell during a pump period, even when the corresponding band is topologically trivial. Our results open the door to studying nonlinearity-induced Thouless pumping of solitons.
Thouless pumping, which plays an important role in understanding the quantum Hall effect, is a phenomenon where quantized transport arises through the slow periodic variation of a Hamiltonian parameter [1, 2]. The transport requires the complete filling of an energy band so that it is dictated by the Chern number of the occupied band with the varying parameter playing the role of an additional momentum. So far, Thouless pumps have been experimentally observed in photonic systems [3, 4, 5], cold atom systems [6, 7, 8, 9, 10] and other systems [11, 12, 13].
Nonlinearity is a natural feature in various systems, including photonic systems [14, 15, 16, 17, 18, 19, 20] and Bose-Einstein condensates (BECs) [21, 22, 23, 24, 25, 26]. It leads to many intriguing phenomena, such as solitons, which are wave packets that travel without changing their shape. Nonlinearity also gives rise to topology-related phenomena, such as topological bulk [27, 28] and edge solitons [29, 30, 31, 32], as well as nonlinearity-induced topological insulators [33, 34]. In particular, it has recently been found that a soliton arising from nonlinearity can exhibit quantized transport when an underlying Hamiltonian parameter is slowly tuned periodically (called nonlinear Thouless pumping)  [35, 36, 37, 38, 39, 40, 41]. Later, a quantized fractional Thouless pumping of solitons is found by considering a soliton bifurcating from multiple bands [42]. The integer and fractional quantized Thouless pumping are now understood as the flow of the instantaneous one-band and maximally localized multi-band Wannier functions [36, 37, 38, 42], respectively. For the former, the displacement of a soliton is determined by the Chern number of the Bloch band of the linear Hamiltonian of the system [36, 37, 38]. If this mechanism always worked, we would not observe the occurrence that a soliton’s displacement is different from this Chern number. In other words, a soliton would not undergo transport when a linear Hamiltonian is topologically trivial.
Here we surprisingly find an anomalous nonlinear Thouless pump that can exhibit the pumping of a soliton across zero, two, or three unit cells over one cycle, while the Chern number of the band, from which the soliton comes, is C=1C=-1. Our result is thus beyond the previous nonlinear Thouless pump where the displacement is identical to the Chern number. We develop a general theory revealing that this anomalous behavior arises from a transition of a soliton between different Wannier functions by passing through an intersite-soliton (or dipole-soliton) state. Leveraging the discovery of the new nonlinear Thouless pump, we construct an alternative nonlinear model, demonstrating the occurrence of quantized Thouless pumping even when the Chern number of the band of the linear Hamiltonian vanishes (in other words, the linear Thouless pumping does not happen in this system over one cycle). To further characterize the topological properties of these anomalous phenomena, we study a spatial supercell containing multiple unit cells to ensure the well localization of a soliton within the supercell. Subsequently, we calculate the Chern number of a modulated Hamiltonian involving effects of the soliton solution, finding it to be identical to the displacements. Finally, we show the emergence of anomalous nonlinear Thouless pumping in a nonlinear continuous model, which can be experimentally realized in cold atom systems.
General theory.—We start by proposing a general theory showing how the anomalous nonlinear Thouless pump can occur. Consider a one-dimensional (1D) non-interacting Hamiltonian $H^{\mathrm{lin}}(\theta)$ with translation symmetry. The Hamiltonian is controlled by a system parameter $\theta$ which varies from 0 to $2\pi$, satisfying $H^{\mathrm{lin}}(0)=H^{\mathrm{lin}}(2\pi)$. The Hamiltonian at each $\theta$ has Bloch states characterized by momentum $k$. We focus on an isolated Bloch band associated with a set of Wannier functions denoted by $W_{l}(j,\theta)$, where $l$ is an integer denoting the lattice vector in 1D, and $j$ is the spatial coordinate. Without loss of generality, we assume that the Chern number of this band with respect to momentum $k$ and parameter $\theta$ is 1-1 (although our theory is applicable to other Chern numbers as well). Consequently, the center-of-mass position (referred to as the Wannier center) of $W_{l}(j,\theta)$ changes by 1-1 as $\theta$ varies from 0 to $2\pi$, assuming that the length of a unit cell is $1$, as shown by the black lines in Fig. 1. Prior research indicates that, in the presence of nonlinearity, soliton solutions that resemble Wannier functions (referred to as onsite solitons) exist, such that the soliton undergoes a displacement equal to the Chern number over one pump period [35, 36, 37, 38]. Therefore, in this case, the soliton experiences a displacement of 1-1 over one cycle.
Figure 1: Schematic illustration of how anomalous nonlinear Thouless pumping arises. The black lines represent the center-of-mass position of Wannier functions with respect to $\theta$. When a soliton follows a black line, normal nonlinear pumping appears. In contrast, if a soliton transitions between Wannier functions by passing through intersite solitons, whose center-of-mass positions are marked by diagonal crosses, anomalous nonlinear pumping occurs. For example, following the blue and red lines in a results in displacements of 0 and 2-2, respectively, while following the green line in b results in a displacement of 3-3.
In addition to the onsite-soliton solution, previous studies have also investigated intersite (or dipole) soliton solutions [18, 19, 20]. We define an ideal intersite soliton as an equal superposition of two neighboring Wannier functions, e.g., $W_{0}(j,\theta)+W_{1}(j,\theta)]/\sqrt{2}$. Note that while the intersite soliton might be reminiscent of the doublon [43], they are completely different. Although intersite solitons themselves have been extensively studied [44, 45, 46, 47, 48, 49, 50, 51, 52], in the context of nonlinear Thouless pumping, prior studies have found that an intersite soliton is always unstable [36], thus precluding the stable evolution between onsite and intersite solitons, as a system parameter is varied.
We now show that a soliton can transition from one Wannier function to another by passing through an intersite-soliton state, leading to the anomalous nonlinear Thouless pumping. Specifically, consider a system where, at $\theta=0$, a soliton solution is $W_{0}(j,0)$, with a center-of-mass position of 0. If the soliton remains in the Wannier state as $\theta$ increases, its center-of-mass will follow the black trajectory, as illustrated in Fig. 1, consistent with previous reports. However, we find that near $\theta=\pi$, stable intersite solitons exist, e.g., $\left[W_{0}(j,\pi)+W_{1}(j,\pi)\right]/\sqrt{2}$ and $\left[W_{0}(j,\pi)+W_{-1}(j,\pi)\right]/\sqrt{2}$, with their center-of-mass positions marked by diagonal crosses in Fig. 1a. Through such an intersite soliton, the soliton can transition from $W_{0}(j,0)$ into either $W_{1}(j,2\pi)$ or $W_{-1}(j,2\pi)$, following the blue or red trajectory and yielding quantized transports of 0 and 2-2, respectively, both of which are distinct from the Chern number. Although the former case is similar to the trapped soliton for strong nonlinearity [35, 37], the mechanism is completely different (see Supplementary Information Section S-1 for details). The trapped phenomenon arises from all band contributions due to strong nonlinearity, leading to zero Chern number [42]. In contrast, in our case, the soliton always occupies one single band.
Moreover, as shown in Fig. 1b, a soliton can first evolve from $W_{0}(j,0)$ to $W_{-1}(j,\pi)$ through an intersite-soliton state formed by $W_{0}$ and $W_{-1}$ and subsequently evolve to $W_{-2}(j,2\pi)$ through an intersite-soliton state formed by $W_{-1}$ and $W_{-2}$. This process results in a displacement of 3-3. These anomalous transport outcomes do not align with the Chern number of the linear band. The theory suggests the potential for various anomalous results; however, it is crucial that the soliton solutions remain stable throughout the entire period.
Figure 2: Anomalous nonlinear Thouless pumping in the discrete nonlinear model. a, One-cycle trajectory of center-of-mass positions of instantaneous solitons coming from the lowest band of the linear Hamiltonian in Eq. (1) as a system parameter $\theta$ varies. It plots three anomalous cases corresponding to displacements of 0 (case 1), 2-2 (case 2), and 3-3 (case 3), described by the blue (g=1g=-1, $g_{12}=0$, and $m_{0}=1$), red ($g=1$, $g_{12}=0$, and $m_{0}=1$), and green lines ($g=1$, $g_{12}=0$, and $m_{0}=1.3$), respectively, in contrast to the normal case with a displacement of 1-1 described by the black line ($g=g_{12}=m_{0}=1$). The associated density distributions $|\psi_{j}|^{2}=\sum_{\sigma}|\psi_{\sigma j}|^{2}$ of the solitons for the three anomalous cases are plotted in b-d. e-h, Occupations $\rho_{W_{l}}(\theta)$ of instantaneous solitons on Wannier functions $W_{l}(\theta)$ of the lowest band along the Wannier centers as a function of $\theta$ for the four cases. See also the Supplementary Information Section S-3 for more information. Here, we set $N=1.45$.
Discrete nonlinear model.—To demonstrate that the anomalous nonlinear Thouless pumping can arise in concrete models, we study the following dimensionless discrete nonlinear Schrödinger equation:
$i\frac{\partial}{\partial t}\psi_{\sigma j}=\sum_{\sigma^{\prime}j^{\prime}}H_{\sigma j,\sigma^{\prime}j^{\prime}}^{\textrm{lin}}(\theta)\psi_{\sigma^{\prime}j^{\prime}}+{V_{\sigma j}(\psi_{\sigma j},\psi_{\overline{\sigma}j})\psi_{\sigma j}},$ (1)
[…]
$H^{\textrm{lin}}(k)=(m_{z}+J_{1}\cos k)\sigma_{z}+J_{1}^{\prime}\sin k\sigma_{y}+J_{2}\sigma_{x},$ (2)
where $\sigma_{\nu}$ ($\nu=x,y,z$) are Pauli matrices, and $m_{z}=m_{0}+\cos\theta$ with $m_{0}$ being a real system parameter. We set $J_{1}=J_{1}^{\prime}=1$ and $J_{2}=\sin\theta$. The parameter $\theta$ is slowly varied as time evolves, realizing a periodic change of the linear Hamiltonian. This model can be transformed to the Rice-Mele model through a unitary transformation realized by $U=e^{-i\pi\sigma_{y}/4}$, and it clearly realizes the Chern band [53] if we view $\theta$ as the other momentum $k_{y}$ besides $k$. The Chern number of the lowest band with respect to $k$ and $\theta$ is $C=1$ when $2<m_{0}<0$, C=1C=-1 when $0<m_{0}<2$, and $C=0$ otherwise. For linear Thouless pumping, particles are required to fill an entire band so that the cloud of particles will travel across the lattice distance identical to the Chern number as we slowly tune $\theta$ from 0 to $2\pi$ [1]. Note that this tight-binding model can be experimentally implemented in ultracold atomic gases as proposed in Ref. [54] with $g$ and $g_{12}$ describing the intraspecies and interspecies interactions, respectively (see the section on nonlinear continuous model).
To demonstrate the anomalous transport behavior in this nonlinear model, we calculate stable instantaneous soliton solutions $\chi_{\sigma j}(\theta)$ that come from the lowest band with the Chern number C=1C=-1 for the instantaneous nonlinear Hamiltonian at each $\theta$ using the Newton’s method; the solution is related to $\psi_{\sigma j}$ through $\psi_{\sigma j}=e^{-i\mu(\theta)t}\chi_{\sigma j}(\theta)$, where $\mu(\theta)$ is the chemical potential. This approach is justified because, in the adiabatic limit—where $\theta$ is varied very slowly—the time evolution of a soliton follows the instantaneous eigenstate of the instantaneous nonlinear Hamiltonian at time $t$ [55, 56, 57, 58, 59]. We have directly computed the time evolution based on Eq. (1) when $\theta$ is varied sufficiently slowly and found that the results closely resemble the instantaneous solutions. Additionally, we have performed a stability analysis to confirm that the solved nonlinear instantaneous soliton solutions remain stable at each $\theta$ (see also Supplementary Information Section S-2).
Figure 3: Transition between normal and anomalous nonlinear Thouless pumping. a-d, Center-of-mass positions of stable instantaneous solitons (black and red lines) and Wannier functions (dashed grey lines) with respect to a system parameter $\theta$ over one cycle for different $g_{12}$. The anomalous nonlinear pumping arises due to appearance of intersite soliton solutions around $\theta=\pi$ (see red lines in b). See Supplementary Information Section S-5 for stability analysis at $\theta=\pi$. Here, we set $m_{0}=g=1$ and $N=1.45$.
Figure 2a illustrates that when $g_{12}=1$, the soliton’s center of mass travels across the lattice by 1-1 over one period, which is equal to the Chern number of the lowest band. This result is consistent with the previous findings [35, 36, 37, 38]. Surprisingly, we observe that when $g_{12}=0$, the soliton’s displacement changes to 0, 2-2, or 3-3, which differ from the Chern number of the lowest band (see also Figs. 2b-d for evolution of the soliton profiles). Here, we set $g=1$ (see Supplementary Information Section S-4 for discussion on effects of $g$ on soliton’s width).
[…]
We proceed to examine how nonlinear Thouless pumping transitions to the anomalous one. To illustrate this, we plot in Fig. 3 the center-of-mass position of stable nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as a function of $\theta$ for different values of $g_{12}$. Notably, when $g_{12}=0.3$, the center-of-mass positions of a soliton and the Wannier centers exhibit clear differences for most values of $\theta$, but they coincide at $\theta=0$ and $\theta=\pi$. In fact, at $\theta=\pi$, the soliton solution is precisely the Wannier function multiplied by $\sqrt{N}$ when $m_{0}=1$ (see Supplementary Information Section S-5). However, as $g_{12}$ decreases, a stable intersite soliton solution emerges at $\theta=\pi$, localized around $j=l$ with $l$ being an integer, differing from the Wannier center, which is located at $l+1/2$. With a further decrease in $g_{12}$, we observe that the new branch of solutions becomes longer, eventually connecting to the original soliton solutions (represented by black lines). Consequently, when $g_{12}<0.2$, a soliton becomes displaced by 2-2 during adiabatic evolution.
Nonlinearity-induced Thouless pumping.—We have demonstrated the breakdown of the correspondence between displacement and the Chern number for a nonlinear Thouless pump, which naturally leads to a question whether a soliton can be pumped when the Chern number of the linear band is zero. Our answer is affirmative. To illustrate this, we again consider the linear Hamiltonian in Eq. (2). Now, we vary the system parameters $m_{z}$, $J_{1}=J_{1}^{\prime}$ and $J_{2}$ with respect to $\theta$. For simplicity, we impose the conditions that all these parameters are symmetric about $\theta=\pi$ as illustrated in Fig. 4a. Consequently, the Hamiltonian $H^{\textrm{lin}}(k,\theta)$ and the Zak phase $\gamma_{n}(\theta)$ also have this symmetry; $\gamma_{n}(\theta)=i\int_{0}^{2\pi}{\rm d}k\langle u_{n,k}|\partial_{k}u_{n,k}\rangle$ is the Zak phase of the 1D system at a fixed $\theta$ with $|u_{n,k}\rangle$ being the $n$th eigenstate of $H^{\textrm{lin}}(k,\theta)$. The Chern number for each band is thus zero so that the linear Thouless pumping cannot occur.
To achieve nonlinearity-induced Thouless pumping of solitons, we simultaneously adjust $g_{12}$, as shown in Fig. 4b (the interaction can be controlled via Feshbach resonances in ultracold atomic gases [25]). Specifically, we initially reduce $g_{12}$ to zero gradually, which facilitates the soliton’s travel over a longer distance. Subsequently, we slowly increase $g_{12}$ back to its original value to prevent the soliton from returning to its initial position. In Fig. 4c, we plot the density distribution of nonlinear eigenstates of the instantaneous nonlinear Hamiltonian as $\theta$ varies from 0 to $2\pi$, remarkably demonstrating that the soliton is displaced by 1-1 unit cell. This effect is further evidenced in Fig. 4d, which depicts the evolution of center-of-mass positions of instantaneous soliton solutions as $\theta$ changes. This observation starkly contrasts with the behavior of the Wannier center (or the Zak phase), which returns to its starting value after one complete cycle. Hence, we successfully achieve Thouless pumping of a soliton through the influence of nonlinearity.
Figure 4: Nonlinearity-induced Thouless pumping. a,b, Illustration of how the parameters $m_{z}(\theta)$, $J_{1}(\theta)$, $J_{2}(\theta)$ and $g_{12}(\theta)$ are varied with respect to $\theta$ in order to induce a nonlinear pumping for a topologically trivial band. c, One-cycle evolution of the density distribution $|\psi_{j}|^{2}$ of instantaneous solitons bifurcating from the lowest band. d, The evolution of center-of-mass positions of instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one cycle. Here, we set $N=1.25$ and $g=1$.
To characterize the topological property of anomalous nonlinear Thouless pumping, we calculate the Chern number of a modulated linear Hamiltonian in a supercell consisting of $L$ unit cells incorporating the effects of solitons [35]. To establish the validity of this method, consider a soliton solution $\psi_{\sigma j}^{\textrm{s}}(\theta)$ that satisfies Eq. (1). This solution is used to define a new linear Hamiltonian at each $\theta$ as
$H^{\prime}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta):=H^{\textrm{lin}}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta)+V_{\sigma j}(\psi_{\sigma j}^{\textrm{s}}(\theta),\psi_{\overline{\sigma}j}^{\textrm{s}}(\theta))\delta_{jj^{\prime}}\delta_{\sigma\sigma^{\prime}},$ (3)
where $V_{\sigma j}(\psi_{\sigma j}^{\textrm{s}}(\theta),\psi_{\overline{\sigma}j}^{\textrm{s}}(\theta))$ is the effective potential generated by the soliton. We then study the time-dependent linear equation, $i\partial_{t}\Psi_{\sigma j}=\sum_{\sigma^{\prime}j^{\prime}}H^{\prime}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta)\Psi_{\sigma^{\prime}j^{\prime}}$, where $H^{\prime}$ is independent of $\Psi$. Clearly, $\psi_{\sigma j}^{\textrm{s}}(\theta)$ is a solution to this linear equation, thus converting a nonlinear problem to a linear one. However, $H^{\prime}$ does not preserve translational symmetry due to the soliton-induced defect potential. To restore the symmetry, we partition the lattice into sections $\dots,[1,L],[L,2L],\dots$, each called a supercell, composed of $L$ unit cells. $L$ is sufficiently large so that a soliton is well localized within a supercell at $\theta=0$. At this $\theta$, a soliton exists within each supercell, as illustrated in Fig. 5a, making the effective potential caused by the solitons periodic with a period of $L$. Since the solitons are far away from each other, the soliton array should also be the solution to Eq. (1). Similar to $H^{\prime}(\theta)$, we define a linear Hamiltonian $H^{\textrm{sc}}(\theta)$ using the effective potential of the soliton array at each $\theta$, which respects the translational symmetry, i.e., $H^{\textrm{sc}}_{\sigma j+L,\sigma^{\prime}j^{\prime}+L}(\theta)=H^{\textrm{sc}}_{\sigma j,\sigma^{\prime}j^{\prime}}(\theta)$.
Our numerical calculations indicate that the soliton in each supercell closely resembles a Wannier function of the Hamiltonian $H^{\textrm{sc}}(\theta)$. This is evidenced by the uniform projections of the soliton wavefunction onto the second band in Fig. 5b, as shown in Fig. 5c, indicating that the soliton wavefunctions are equal superpositions of all the Bloch states in this band, and are thus Wannier functions. Since the displacement of a Wannier function over a pump period is determined by the Chern number $C_{\textrm{sc}}$ of this band of $H^{\textrm{sc}}$, the soliton’s displacement is thus dictated by this Chern number. To see this, we introduce a new parameter $\theta^{\prime}=\theta/L$ such that $H^{\textrm{sc}}(\theta^{\prime}+2\pi)=H^{\textrm{sc}}(\theta^{\prime})$. As $\theta^{\prime}$ varies from 0 to $2\pi$, the Wannier function travels by $C_{\textrm{sc}}$ supercells, equivalent to $C_{\textrm{sc}}L$ original unit cells, indicating that the average displacement over a cycle of $\theta$ is $C_{\textrm{sc}}$. Our calculations reveal that for normal nonlinear pumping, $C_{\textrm{sc}}$ is equal to the Chern number $C$ of the corresponding Bloch band of $H^{\textrm{lin}}$; however, they differ in the anomalous one. For example, in the anomalous cases shown in Fig. 2, $C_{\textrm{sc}}$ can be 0, 2-2, or 3-3 , while C=1C=-1. In Fig. 4, $C_{\textrm{sc}}=-1$ while $C=0$. In addition, we find that $C_{\textrm{sc}}$ changes from 1-1 to 2-2 as $g_{12}$ decreases across $0.2$ in Fig. 3.
Figure 5: Supercell method. a, The effective potential generated by a soliton array with one soliton per supercell (three supercells are explicitly shown, each containing $L=40$ sites) as $\theta^{\prime}=\theta/L$ varies from 0 to $2\pi$. The potential is periodic with period $L$. Only the potential felt by the first component is plotted, and the potential for the other component is similar. b, Energy spectra of the modulated linear Hamiltonian $H^{\textrm{sc}}$ containing the linear potentials in a with respect to momentum $k$ and $\theta^{\prime}$. Given that the spectra are periodic functions of $\theta^{\prime}$ with a period of $2\pi/L$, we only plot part of the spectrum with $\theta^{\prime}\in[0,2\pi/L]$. Here, we show three Bloch bands around $\mu(\theta)$ and the second band corresponds to the soliton wavefunction $\psi_{\sigma j}^{\textrm{s}}(\theta)$. c, The normalized projection of a soliton wavefunction $\psi_{\sigma j}^{\textrm{s}}$ on the second band in b, $\rho_{\textrm{sc}}(k,\theta^{\prime})=\frac{L_{\textrm{sc}}}{N}|\sum_{\sigma j}\varphi_{k,\sigma,j}(\theta^{\prime})^{*}\psi_{\sigma j}^{\textrm{s}}(\theta^{\prime})|^{2},$ where $L_{\textrm{sc}}$ is the number of supercells, $\varphi_{k,\sigma,j}(\theta^{\prime})$ is a Bloch state at momentum $k\in[0,2\pi/L]$ in the corresponding band of the Hamiltonian $H^{\textrm{sc}}$. Here, we consider case 2 in Fig. 2 with $N=1.45$, $m_{0}=1$, $g_{12}=0$, and $g=1$.
Continuous nonlinear model.—To implement the tight-binding model in Eq. (2) in ultracold atomic gases, we consider the following continuous model:
$H_{\textrm{c}}^{\textrm{lin}}{(x,\theta)}=H_{0}{(x)}+h_{z}{(\theta)}\sigma_{z}+V_{\textrm{so}}{(x,\theta)}\sigma_{x},$ (4)
where the spin is encoded in two hyperfine states of an atom. Note that similar models have been proposed for realizing linear Thouless pumping in both two-dimensional and three-dimensional cold atom systems [54]. In this Hamiltonian, $H_{0}{(x)}=\frac{p_{x}^{2}}{2m}-V_{x}\cos^{2}(k_{R}x)$, where $p_{x}=-i\hbar\partial_{x}$ is the momentum operator, $m$ is the mass of atoms, and $V_{x}$ is the strength of optical lattices generated by lasers with wavevector $k_{R}$. In addition, $h_{z}$ denotes the strength of the Zeeman field corresponding to the detuning of the Raman lasers, and $V_{\textrm{so}}{(x,\theta)}=V_{s}\sin(k_{R}x)+V_{c}\cos(k_{R}x)\sin(\theta)$. This potential can be implemented using two pairs of Raman lasers, as described in Ref. [54, 60, 61, 62]. One pair (the other) contains a laser beam with the Rabi frequency proportional to $\sin(k_{R}x)$ $\cos(k_{R}x)$] and another beam proportional to $e^{ik_{R}y}$. Given that our system is effectively 1D due to a strong confinement along $y$ and $z$, the effects of the phase $e^{ik_{R}y}$ can be negligible. The Hamiltonian $H_{\textrm{c}}^{\textrm{lin}}(x,\theta)$ is invariant under translation by $2a$ with $a=\pi/k_{R}$. Despite this, we show that its Bloch states can still be characterized by the momentum $k$ within the interval $0,2\pi/a]$, due to the symmetry represented by $T_{a}\sigma_{z}$, where $T_{a}$ is the translation operator by $a$ (see Supplementary Information Section S-6). As a result, for the linear Thouless pumping in this model, the average displacement per particle over one pump period is $x_{c}=Ca$, where $C$ is the Chern number for the corresponding Bloch band (see Supplementary Information Section S-6).
The continuous model can be mapped to a tight-binding model based on the Wannier functions $\{W_{nj}(x):j\in\mathbb{Z}\}$ from the lowest Bloch band of the Hamiltonian $H_{0}(x)$ with the band index $n=1$. This yields precisely the Hamiltonian in Eq. (2) with the parameters $J_{1}=-2\int{\rm d}xW_{10}(x)H_{0}(x)W_{11}(x)$, $J_{1}^{\prime}=-2V_{s}\int{\rm d}xW_{10}(x)\sin(k_{R}x)W_{11}(x)$, $J_{2}=V_{c}\sin\theta\int{\rm d}xW_{10}(x)\cos(k_{R}x)W_{10}(x)$, and $m_{z}=h_{z}$ (see the derivation in Supplementary Information Section S-6). Thus, we expect the emergence of anomalous nonlinear Thouless pumping in the continuous model in the presence of interactions described by nonlinear terms. In our study, we set $V_{x}=4E_{R}$, $V_{s}=1.04E_{R}$ and $V_{c}=0.21E_{R}$, where $E_{R}=\hbar^{2}k_{R}^{2}/(2m)$ is the recoil energy. The parameter $h_{z}$ is varied as $h_{z}=2t_{x}(1+\cos\theta)$ with $t_{x}=0.0855E_{R}$. Under these conditions, we find that the Chern number of the first and second bands of the continuous model are $C=1$ and C=1C=-1, respectively. As a result, the average displacement for the linear Thouless pumping over one cycle is $a$ and a-a, respectively (see Fig. 6b).
We now demonstrate the emergence of anomalous nonlinear pumping in the continuous model in the presence of nonlinearity. The dynamics is governed by the following dimensionless Gross-Pitaevskii (GP) equation:
$i\frac{\partial}{\partial t}\psi_{\sigma}(x)=[\tilde{H}_{c}^{\textrm{lin}}\psi(x)]_{\sigma}+g|\psi_{\sigma}(x)|^{2}\psi_{\sigma}(x),$ (5)
where $\psi=\left(\psi_{1}(x),\psi_{2}(x)\right)^{T}$ and $\tilde{H}_{c}^{\textrm{lin}}$ is the dimensionless version of ${H}_{c}^{\textrm{lin}}$. The units of energy, time and length are $E_{R}$, $\hbar/E_{R}$ and $a$, respectively. We have taken $g_{11}=g_{22}=g$ for simplicity and have confirmed that the results remain valid when they are slightly different. We set $g=1$ and g=1g=-1 for repulsive and attractive interactions, respectively. The strength of nonlinearity is characterized by the norm $N=\sum_{\sigma}\int dx|\psi_{\sigma}(x)|^{2}$. For instance, consider a $^{7}\text{Li}$ BEC containing $1316$ atoms with the scattering length $a_{s}\approx-1.43$ nm under optical lattices along $x$ with $a=532$ nm and a transverse harmonic trap along $y$ and $z$ with the frequency $\omega_{\perp}=2\pi\times 710$ Hz [63, 64]. In this case, we have $N=0.2$.
By solving the nonlinear eigenstates of the instantaneous nonlinear Hamiltonian, we find that when $g>0$, the displacement of a soliton bifurcating from the first band over one period is $a$, consistent with linear Thouless pumping. However, the corresponding tight-binding model produces a nonlinear pumping displacement of $2a$, suggesting that the tight-binding model may oversimplify the nonlinear effects (see the discussion in Supplementary Information Section S-6). Remarkably, we find that when $g<0$, a soliton that bifurcates from the second band in the continuous model is pumped across 2a-2a over one cycle, as depicted in Fig. 6, which is twice the Chern number of the second band. We also find that for negative $g$, a soliton coming from the first band (whose Chern number is $1$) does not exhibit any pumping over one cycle (see Supplementary Information Section S-1), corresponding to the first anomalous case shown in Fig. 1a and Fig. 2b. Furthermore, we have numerically verified the stability of the instantaneous soliton solutions as $\theta$ runs from 0 to $2\pi$. In addition, the numerical time evolution of an initial soliton is also found to be in agreement with the instantaneous solution. These results suggest that anomalous nonlinear pumping can be experimentally observed in the continuous model.
Figure 6: Anomalous nonlinear pumping in the continuous model. a, One-cycle evolution of the density distribution $|\psi(x)|^{2}=\sum_{\sigma}|\psi_{\sigma}(x)|^{2}$ of the instantaneous solitons bifurcating from the second band when $g_{12}=0$. Note that the anomalous nonlinear pumping can still occur when $g_{12}\neq 0$. b, The evolution of center-of-mass positions $x_{c}$ of the instantaneous solitons (red line) and the corresponding Wannier functions (black line) with respect to $\theta$ over one period. Here, N=0.2{N}=0.2 and g=1g=-1.
Discussion.—We have demonstrated the emergence of anomalous nonlinear Thouless pumping, wherein the correspondence between the displacement of a soliton and the Chern number of the underlying linear Bloch band breaks down. This anomalous behavior surpasses the understanding of nonlinear Thouless pumping based on the center-of-mass flow of the instantaneous Wannier functions. We show that the breakdown arises due to the transition of a soliton between different Wannier functions by passing through an intersite-soliton state. Furthermore, we illustrate that nonlinearity can induce the Thouless pumping of a soliton, even when the corresponding linear Bloch band is topologically trivial. We note that our results are completely different from the phenomenon in a Bose-Bose atomic mixture in Ref. [38] where a soliton of impurity atoms can undergo a nonzero quantized displacement during a pump cycle, as a result of the dragging from a soliton of majority atoms that bifurcates from a topologically nontrivial band. There, the pumping can also be understood as the center-of-mass flow of the Wannier functions of the corresponding Bloch band of the linear Hamiltonian felt by the majority atoms. In our case, we have only one set of band structures, and the discrepancy between the band topology and quantized motion emerges spontaneously, going beyond the description of the Wannier function of the original linear Hamiltonian. Given that anomalous nonlinear Thouless pumping can occur in both discrete and continuous models, it can be experimentally observed in either photonic systems or ultracold atomic gases. Additionally, our work inspires the exploration of other interesting topics, such as anomalous fractional Thouless pumping of solitons.
Acknowledgments We thank Y. Zhang and Y.-B. Yang for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604). We also acknowledge the support by center of high performance computing, Tsinghua University.